A string is 6.17 m long with a mass of 3.37 g. The string is held taut with a tension of 448 N applied to the string. A pulse is sent down the string. How long does it take the pulse to travel the 3.00 m of the string?




Round your answer to 5 decimal places.



Do not forget the unit

1
Expert's answer
2022-05-11T13:53:07-0400
"v=\sqrt{F/(m/l)}"

"v=\sqrt{448/(0.00337/6.17)}=906\:\rm m/s"

"\Delta t=\frac{d}{v}=\frac{3.00}{906}=0.0033\:\rm s"